# Class-Aware PillarMix: Can Mixed Sample Data Augmentation Enhance 3D Object Detection with Radar Point Clouds?

### v2

@@ IV Experiments
TABLE I: Comparison of $AP_{BEV}$/$AP_{3D}$ of different MSDA methods on Bosch Street dataset.
@@
CutMix [14] 66.05/56.71 35.98/32.40 11.82/7.70 15.83/11.79 32.42/27.15 70.78/61.42 44.04/39.99 15.77/10.31 19.53/15.26 37.53/31.78
MixUp [11] 66.11/56.81 39.01/34.89 15.46/9.81 17.81/13.48 34.60/28.73 70.73/61.94 46.13/42.32 17.11/11.12 21.50/16.82 38.86/32.93
LaserMix [2] 66.05/55.95 38.43/33.84 14.85/9.22 17.73/13.61 34.27/28.16 70.81/61.42 46.36/42.50 16.49/10.63 20.63/16.24 38.57/32.70
PolarMix [4] 65.27/56.25 35.40/31.23 13.42/8.83 16.84/12.36 32.73/27.17 69.97/60.73 44.28/40.21 16.03/10.18 20.28/16.11 37.64/31.81
PillarMix [5] 64.87/55.73 40.22/35.89 15.58/9.98 15.78/11.47 34.11/28.27 70.23/61.17 48.15/44.54 16.70/10.92 19.90/15.56 38.75/33.05
@@
CutMix [14] 74.39/66.30 49.20/44.95 17.75/11.96 23.11/18.55 41.11/35.44 76.47/69.11 53.22/49.21 20.38/14.57 26.44/22.17 44.13/38.76
MixUp [11] 74.57/66.44 51.51/47.59 19.27/13.25 25.07/20.47 42.60/36.93 76.49/69.37 55.80/52.05 21.14/15.27 28.28/23.70 45.43/40.10
LaserMix [2] 74.53/65.48 50.52/46.59 18.69/12.22 24.62/19.73 42.08/35.93 76.25/67.77 54.72/50.75 20.46/13.92 28.24/23.08 44.92/38.88
PolarMix [4] 73.97/66.03 47.77/44.38 18.10/12.28 23.47/19.16 40.82/35.46 74.89/67.67 52.80/48.45 20.10/14.40 27.34/22.62 43.78/39.29
PillarMix [5] 74.23/66.00 52.38/48.47 18.89/13.15 23.63/18.83 42.28/36.61 76.15/68.89 55.50/51.68 20.57/14.79 27.15/22.26 44.84/39.30
@@
[…]
@@ IV-A Datasets
    Bosch Street [12] is a large-scale, multi-modal dataset featuring imaging radar. It comprises 1.3 million frames (36.5 hours) collected across nine cities under three weather conditions. We extract a subset by subsampling at 1-second intervals, resulting in 81k frames for training and 24k for validation, which serve as the $100\%$ data baseline in subsequent experiments. We conduct experiments on datasets of different scales, where the subsets are generated by randomly selecting entire sequences rather than individual frames, reflecting a practical data collection workflow. All results are reported on the validation set. To obtain higher-quality point clouds, we apply a five-frame aggregation strategy. For more details, we refer readers to the original paper [12].
@@
    K-Radar [13] contains over 35k frames of around one hour recorded in two cities under seven weather conditions. As noted previously, in the original dataset split, frames in the training and test sets were sampled from the same sequences [20, 21]. To address this, we restructured the dataset by sequences, assigning 44 sequences (27k frames) to the training set and the remaining 24 sequences (7.9k frames) to the test set. The new split ensures that all weather conditions are present in both the training and test sets. Since the number of sequences is limited, the subsets for experiments are randomly sampled by frames. For runtime efficiency, we generate the sparse point cloud using constant false alarm rate (CFAR) thresholding on the raw 4D radar cube, rather than deriving points only from the top 10% of power values as in the original paper, which aligns with standard radar signal processing practices [18].
@@ IV-B Implementation Details
For evaluation, we adopt CenterPoint [22] as our 3D detector. The detection range for the $x$, $y$, and $z$ axes is defined as $[0,80]\times[-16,16]\times[-2,7.6]$ meters for both datasets. Following K-Radar [13] settings, we utilize the AdamW optimizer along with a Cosine Annealing learning rate scheduler and evaluate performance using $AP_{BEV}/AP_{3D}$ with an IoU threshold of 0.3 for all classes. Unlike previous setups, we report $AP|_{R_{40}}$ instead of $AP|_{R_{11}}$ as it offers a more representative evaluation [23].
Moving on to the hyperparameter settings of CAPMix, we set the pillar size $\phi=2$ meters (i.e., covering an area of 4 square meters in BEV) for all pillar-based MSDA. As for the choice of class-aware $\beta$ distribution, we use ${\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}}=\beta(2,2)$, ${\color[rgb]{0.4660,0.6740,0.1880}\definecolor[named]{pgfstrokecolor}{rgb}{0.4660,0.6740,0.1880}\beta_{s}}=\beta(4,3)$, ${\color[rgb]{0.6350,0.0780,0.1840}\definecolor[named]{pgfstrokecolor}{rgb}{0.6350,0.0780,0.1840}\beta_{d}}=\beta(0.1,5)$ based on empirical experiments (see Sec. V-C). For the Bosch Street dataset, four classes are partitioned as follows: Car $\rightarrow{\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}}$, Large Vehicle $\rightarrow{\color[rgb]{0.6350,0.0780,0.1840}\definecolor[named]{pgfstrokecolor}{rgb}{0.6350,0.0780,0.1840}\beta_{d}}$, Pedestrian, Ridable Vehicle $\rightarrow{\color[rgb]{0.4660,0.6740,0.1880}\definecolor[named]{pgfstrokecolor}{rgb}{0.4660,0.6740,0.1880}\beta_{s}}$. For K-Radar, only two classes are considered: Sedan $\rightarrow{\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}}$, Bus/Truck $\rightarrow{\color[rgb]{0.6350,0.0780,0.1840}\definecolor[named]{pgfstrokecolor}{rgb}{0.6350,0.0780,0.1840}\beta_{d}}$. To ensure a fair comparison, we follow the procedure outlined in PillarMix [5] by applying random rotations and scaling to all methods, including baseline without MSDA. All methods are implemented using the MMDet3D framework [24]. The reported results represent the mean of three runs with different random seeds on NVIDIA A100 GPU.
@@ V-A Bosch Street Dataset
Table I presents a comparison of different MSDA methods applied to 5%, 10%, 25%, and 50% of the training data. To understand the effectiveness of these methods, we first examine their impact across different object classes. Although MSDA techniques generally improve performance, the degree of enhancement varies. In particular, the car class shows relatively minor gains, likely due to its extensive number of annotations. Even with just 5% of the data, there are still 22,334 annotated cars, compared to only 1,611 large vehicles, 6,495 pedestrians, and 1,816 ridable vehicles. Consequently, an ill-suited MSDA method can actually reduce performance, as seen in 5% dataset results, where LaserMix [2] and PillarMix [5] drop performance by 0.82 and 1.22 on $AP_{BEV}$, respectively.
@@ V-B K-Radar
We extend our study to K-Radar, which focuses on two vehicle classes (sedan and bus/truck) as considered in the original paper [13]. Since K-Radar excludes categories like pedestrian and ridable vehicle, we apply only the ${\color[rgb]{0.6350,0.0780,0.1840}\definecolor[named]{pgfstrokecolor}{rgb}{0.6350,0.0780,0.1840}\beta_{d}}$ and ${\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}}$ components of CAPMix, leaving its advantage for sparse classes untested. Furthermore, K-Radar’s adverse weather conditions introduce additional challenges beyond those found in Bosch Street data.
@@
CutMix [14] 47.60/37.53 24.83/15.86 36.21/26.69
MixUp [11] 49.28/38.30 30.65/23.86 39.97/31.09
LaserMix [2] 46.59/36.52 31.99/25.87 39.29/30.35
PolarMix [4] 47.02/35.62 24.41/14.67 35.71/25.14
PillarMix [5] 47.89/36.96 34.55/28.38 41.22/32.67
@@
[…]
TABLE II: Comparison of $AP_{BEV}$/$AP_{3D}$ of different MSDA methods on K-Radar.
@@ V-C Ablation Study
    Pillar Size Effect: In this part, we examine the impact of pillar size on perception performance. PillarMix [5] has shown that pillar size is crucial, since larger sizes require fewer computations but reduce the effectiveness of edge-truncated augmentation. As shown in Table III, using 8-meter pillars lowers performance across all four classes, with large vehicles experiencing about a 1.4% drop. This drop can be attributed to larger objects spanning multiple pillars, reducing the total number of distinct pillars and limiting augmentation diversity. It is important to note that, unlike PillarMix (which directly swaps pillars between two frames), we apply MixUp within each pillar. This approach selects random pillars from the entire sample instead of relying on a cross-selection pattern, resulting in more diverse augmented point clouds and enhancing the effectiveness of augmentation.

### v3

@@ IV Experiments
TABLE I: Comparison of $AP_{BEV}$$AP_{3D}$ of different MSDA methods on Bosch Street dataset.
@@
CutMix [yun2019cutmix] 66.05/56.71 35.98/32.40 11.82/7.70 15.83/11.79 32.42/27.15 70.78/61.42 44.04/39.99 15.77/10.31 19.53/15.26 37.53/31.78
MixUp [zhang2017mixup] 66.11/56.81 39.01/34.89 15.46/9.81 17.81/13.48 34.60/28.73 70.73/61.94 46.13/42.32 17.11/11.12 21.50/16.82 38.86/32.93
LaserMix [kong2023lasermix] 66.05/55.95 38.43/33.84 14.85/9.22 17.73/13.61 34.27/28.16 70.81/61.42 46.36/42.50 16.49/10.63 20.63/16.24 38.57/32.70
PolarMix [xiao2022polarmix] 65.27/56.25 35.40/31.23 13.42/8.83 16.84/12.36 32.73/27.17 69.97/60.73 44.28/40.21 16.03/10.18 20.28/16.11 37.64/31.81
PillarMix [wu2024pillarmix] 64.87/55.73 40.22/35.89 15.58/9.98 15.78/11.47 34.11/28.27 70.23/61.17 48.15/44.54 16.70/10.92 19.90/15.56 38.75/33.05
@@
CutMix [yun2019cutmix] 74.39/66.30 49.20/44.95 17.75/11.96 23.11/18.55 41.11/35.44 76.47/69.11 53.22/49.21 20.38/14.57 26.44/22.17 44.13/38.76
MixUp [zhang2017mixup] 74.57/66.44 51.51/47.59 19.27/13.25 25.07/20.47 42.60/36.93 76.49/69.37 55.80/52.05 21.14/15.27 28.28/23.70 45.43/40.10
LaserMix [kong2023lasermix] 74.53/65.48 50.52/46.59 18.69/12.22 24.62/19.73 42.08/35.93 76.25/67.77 54.72/50.75 20.46/13.92 28.24/23.08 44.92/38.88
PolarMix [xiao2022polarmix] 73.97/66.03 47.77/44.38 18.10/12.28 23.47/19.16 40.82/35.46 74.89/67.67 52.80/48.45 20.10/14.40 27.34/22.62 43.78/39.29
PillarMix [wu2024pillarmix] 74.23/66.00 52.38/48.47 18.89/13.15 23.63/18.83 42.28/36.61 76.15/68.89 55.50/51.68 20.57/14.79 27.15/22.26 44.84/39.30
@@
  • The best and second best scores are highlighted in bold and underline. The symbol $\triangle\uparrow$ indicates the increase compared to the baseline without MSDA.
@@ IV-A Datasets
    Bosch Street [armanious2024bosch] is a large-scale, multi-modal dataset featuring imaging radar. It comprises 1.3 million frames (36.5 hours) collected across nine cities under three weather conditions. We extract a subset by subsampling at 1-second intervals, resulting in 81k frames for training and 24k for validation, which serve as the $100\%$ data baseline in subsequent experiments. We conduct experiments on datasets of different scales, where the subsets are generated by randomly selecting entire sequences rather than individual frames, reflecting a practical data collection workflow. All results are reported on the validation set. To obtain higher-quality point clouds, we apply a five-frame aggregation strategy. For more details, we refer readers to the original paper [armanious2024bosch].
@@
    K-Radar [paek2022k] contains over 35k frames of around one hour recorded in two cities under seven weather conditions. As noted previously, in the original dataset split, frames in the training and test sets were sampled from the same sequences [fent2024dpft, zhang2024exploring]. To address this, we restructured the dataset by sequences, assigning 44 sequences (27k frames) to the training set and the remaining 24 sequences (7.9k frames) to the test set. The new split ensures that all weather conditions are present in both the training and test sets. Since the number of sequences is limited, the subsets for experiments are randomly sampled by frames. For runtime efficiency, we generate the sparse point cloud using constant false alarm rate (CFAR) thresholding on the raw 4D radar cube, rather than deriving points only from the top 10% of power values as in the original paper, which aligns with standard radar signal processing practices [han2023radar].
@@ IV-B Implementation Details
For evaluation, we adopt CenterPoint [yin2021center] as our 3D detector. The detection range for the $x$, $y$, and $z$ axes is defined as $0,80]\times[-16,16]\times[-2,7.6]$ meters for both datasets. Following K-Radar [paek2022k] settings, we utilize the AdamW optimizer along with a Cosine Annealing learning rate scheduler and evaluate performance using $AP_{BEV}/AP_{3D}$ with an IoU threshold of 0.3 for all classes. Unlike previous setups, we report $AP|_{R_{40}}$ instead of $AP|_{R_{11}}$ as it offers a more representative evaluation [simonelli2019disentangling].
Moving on to the hyperparameter settings of CAPMix, we set the pillar size $\phi=2$ meters (i.e., covering an area of 4 square meters in BEV) for all pillar-based MSDA. As for the choice of class-aware $\beta$ distribution, we use ${\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}}=\beta(2,2)$, ${\color[rgb]{0.4660,0.6740,0.1880}\definecolor[named]{pgfstrokecolor}{rgb}{0.4660,0.6740,0.1880}\beta_{s}}=\beta(4,3)$, ${\color[rgb]{0.6350,0.0780,0.1840}\definecolor[named]{pgfstrokecolor}{rgb}{0.6350,0.0780,0.1840}\beta_{d}}=\beta(0.1,5)$ based on empirical experiments (see Sec. V-C). For the Bosch Street dataset, four classes are partitioned as follows: Car $\rightarrow{\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}}$, Large Vehicle $\rightarrow{\color[rgb]{0.6350,0.0780,0.1840}\definecolor[named]{pgfstrokecolor}{rgb}{0.6350,0.0780,0.1840}\beta_{d}}$, Pedestrian, Ridable Vehicle $\rightarrow{\color[rgb]{0.4660,0.6740,0.1880}\definecolor[named]{pgfstrokecolor}{rgb}{0.4660,0.6740,0.1880}\beta_{s}}$. For K-Radar, only two classes are considered: Sedan $\rightarrow{\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}}$, Bus/Truck $\rightarrow{\color[rgb]{0.6350,0.0780,0.1840}\definecolor[named]{pgfstrokecolor}{rgb}{0.6350,0.0780,0.1840}\beta_{d}}$. To ensure a fair comparison, we follow the procedure outlined in PillarMix [wu2024pillarmix] by applying random rotations and scaling to all methods, including baseline without MSDA. All methods are implemented using the MMDet3D framework [mmdet3d2020]. The reported results represent the mean of three runs with different random seeds on NVIDIA A100 GPU.
@@ V-A Bosch Street Dataset
Table I presents a comparison of different MSDA methods applied to 5%, 10%, 25%, and 50% of the training data. To understand the effectiveness of these methods, we first examine their impact across different object classes. Although MSDA techniques generally improve performance, the degree of enhancement varies. In particular, the car class shows relatively minor gains, likely due to its extensive number of annotations. Even with just 5% of the data, there are still 22,334 annotated cars, compared to only 1,611 large vehicles, 6,495 pedestrians, and 1,816 ridable vehicles. Consequently, an ill-suited MSDA method can actually reduce performance, as seen in 5% dataset results, where LaserMix [kong2023lasermix] and PillarMix [wu2024pillarmix] drop performance by 0.82 and 1.22 on $AP_{BEV}$, respectively.
@@ V-B K-Radar
We extend our study to K-Radar, which focuses on two vehicle classes (sedan and bus/truck) as considered in the original paper [paek2022k]. Since K-Radar excludes categories like pedestrian and ridable vehicle, we apply only the $\color[rgb]{0.6350,0.0780,0.1840}\definecolor[named]{pgfstrokecolor}{rgb}{0.6350,0.0780,0.1840}\beta_{d}$} and $\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}$} components of CAPMix, leaving its advantage for sparse classes untested. Furthermore, K-Radar’s adverse weather conditions introduce additional challenges beyond those found in Bosch Street data.
@@
CutMix [yun2019cutmix] 47.60/37.53 24.83/15.86 36.21/26.69
MixUp [zhang2017mixup] 49.28/38.30 30.65/23.86 39.97/31.09
LaserMix [kong2023lasermix] 46.59/36.52 31.99/25.87 39.29/30.35
PolarMix [xiao2022polarmix] 47.02/35.62 24.41/14.67 35.71/25.14
PillarMix [wu2024pillarmix] 47.89/36.96 34.55/28.38 41.22/32.67
@@
  • The best and second best scores are highlighted in bold and underline. The symbol $\triangle\uparrow$ indicates the increase compared to the baseline without MSDA.
TABLE II: Comparison of $AP_{BEV}$$AP_{3D}$ of different MSDA methods on K-Radar.
@@ V-C Ablation Study
    Pillar Size Effect: In this part, we examine the impact of pillar size on perception performance. PillarMix [wu2024pillarmix] has shown that pillar size is crucial, since larger sizes require fewer computations but reduce the effectiveness of edge-truncated augmentation. As shown in Table III, using 8-meter pillars lowers performance across all four classes, with large vehicles experiencing about a 1.4% drop. This drop can be attributed to larger objects spanning multiple pillars, reducing the total number of distinct pillars and limiting augmentation diversity. It is important to note that, unlike PillarMix (which directly swaps pillars between two frames), we apply MixUp within each pillar. This approach selects random pillars from the entire sample instead of relying on a cross-selection pattern, resulting in more diverse augmented point clouds and enhancing the effectiveness of augmentation.